A jet flew an average speed of 480 mph from point x to point y. Because of headwinds the jet averaged only 460 mph on the return trip. The return trip took 30 mins longer. How many hours was the flight from point y to point x.
Answer by mananth(16946)

let the distance be x
time from x to y = x/480
time from y to x = x/460
time difference = 1/2 hour
x/460 - x/480 = 1/2
(480x-460x)/(460*480) = 1/2
20x = 460*480/2
20x = 460*240
x= 460*240/20
x= 460*12 this is the distance
speed return = 460
time = 460*12/ 460
time = 12 hours on return trip